OpenStudy (anonymous):

Which terms in the expression are like terms? –b – bc + 0.1b – 6b Choose all answers that are correct. -b -bc 0.1b -6b

OpenStudy (zpupster):

the b's are like terms and the bc term is different

OpenStudy (anonymous):

so all of them

sammixboo (sammixboo):

Like terms are numbers with variables or just variables that contain the same variable

OpenStudy (anonymous):

1 2 and 4

OpenStudy (anonymous):

hello? is anyone there i think my screen is frozen

sammixboo (sammixboo):

Not B. there is only one bc.

OpenStudy (texaschic101):

they have to be exactly the same letters... example : a + 4a --- like terms b + ab -- unlike terms ab + ba -- like terms

OpenStudy (anonymous):

ok so all except b

sammixboo (sammixboo):

Correct!

OpenStudy (anonymous):

2 more questions

sammixboo (sammixboo):

Okie dokie

OpenStudy (anonymous):

3(x-5) = 8x -32 X = ?

OpenStudy (anonymous):

-5, 5, -3, or 3

sammixboo (sammixboo):

OK so first distribute the left side of the equation 3(x-5) = 8x -32 3x - 15 = 8x - 32

OpenStudy (anonymous):

????

OpenStudy (anonymous):

5

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

Hello?

sammixboo (sammixboo):

I am confused. It is none of them

OpenStudy (texaschic101):

3(x - 5) = 8x - 32 3x - 15 = 8x - 32 3x - 8x = -32 + 15 -5x = - 17 x = 17/5 or 3 2/5

OpenStudy (texaschic101):

none of your answer choices...was the question written down correctly ?